OpenStudy (anonymous):

Simplify the expression. A. 8 B. 10 C. 16 D. 22

OpenStudy (anonymous):

What is the expression?

OpenStudy (anonymous):

8+3x4 2

OpenStudy (anonymous):

\[\frac{ 8+3\times 4 }{ 2 }\] is this right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So using order of operations (PEMDAS) your first step is to multiply 3*4

OpenStudy (anonymous):

you are then left with \[\frac{ 8+12 }{ 2 }\]

OpenStudy (anonymous):

we are working on the top side first and the bottom after so the only other operation available is addition, can you get the answer now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{ 20 }{ 2 }\]

OpenStudy (anonymous):

22

OpenStudy (anonymous):

8 + 3 x 4 divided by 2 ? 3 * 4 = 12 + 8 = 20 / 2 = 10 the answer is B.

OpenStudy (anonymous):

if you have 20 coins and want to divide between you and your brother how many do each get? \[\frac{ 20 }{ 2 }=10\]

OpenStudy (anonymous):

10

OpenStudy (anonymous):

you anwser more

OpenStudy (anonymous):

I'll try to help

OpenStudy (anonymous):

Simplify the expression. 4(6 – 3) A. 7 B. 12 C. 21 D. 24

OpenStudy (anonymous):

4(6-3)

OpenStudy (anonymous):

6-3 first i think

OpenStudy (anonymous):

post it as a new question

OpenStudy (anonymous):

the expression is 4(6-3)

OpenStudy (anonymous):

close this question and post it as a new question on a separate page

OpenStudy (anonymous):

why

OpenStudy (anonymous):

to get a medal please

OpenStudy (anonymous):

4(6 – 3) 6 - 3 = 3 4 * 3 = 12